OpenStudy (anonymous):

Solve by Substiution 3+4(2y-9)+5x-2y=8 and 3(x+3)-4(2y+1)-2x=4

OpenStudy (anonymous):

http://www.purplemath.com/modules/systlin4.htm this can help you out

OpenStudy (anonymous):

3+8y-36+5x-2y=8 -41+6y+5x=0 -6y=5x-41 y=-6(5x-41) plug in y to other equation

OpenStudy (anonymous):

should get x=7 y=1